Using a long rod that has length , you are going to lay out a square plot in which the length of each 2. side is . Thus the area of the plot will be However, you do not know the value of , so you decide to make n independent measurements X1;X2; :::;Xn of the length. Assume that each Xi has mean 2. (unbiased measurements) and variance1. A large insurance agency services a number of customers who have purchased both a homeowners policy and an automobile pol
Answer #1

Let, X be the deductible amount on the auto policy & Y be the deductible amount on the home owner's policy.

Probability distibution table -

0 100 200 P(x)
100 0.20 0.10 0.20 0.50
250 0.05 0.15 0.30 0.50
P(y) 0.25 0.25 0.50 1

a) -

Probability that $100 deductible on both the policies - P(x=$100, y=$100) =

P(x=$100, y=$100) = 0.10

Hence, probability that $100 deductible on both the policies is 0.10.

b) -

Probability of a deductible amount \geq 100 for homeowner's policy - P(y \geq 100) =

P(y \geq 100) = P(y = 100) + P(y = 200) = 0.25 + 0.50 = 0.75

Probability of a deductible amount \geq 100 for homeowner's policy is 0.75.

c) -

Probability that $100 deductible on auto policy given a $100 deductible on homeowner's policy - P(X=100|Y=100) =

By conditional probability, P(A|B) = \frac{P(A\bigcap B)}{P(B)}

So,

P(X=100|Y=100) = \frac{P(X=100,Y=100)}{P(Y=100)} = \frac{0.10}{0.25} = 0.4

Probability that $100 deductible on auto policy given a $100 deductible on homeowner's policy is 0.4.

d) -

Formula for Cov(X,Y) is as follows -

Cov(X,Y) = E(XY) - E(X)E(Y)

Where,

E(X) = \sum xP(x)

E(X) = 100(0.5) + 250(0.5) = 50 + 125 = 175

E(Y) = \sum yP(y)

E(Y) = 0(0.25) + 100(0.25) + 200(0.5) = 0 + 25 + 100 = 125

E(XY) = \sum_{ij} x_{i}y_{j}P(x,y)

E(XY) = (100)(0)(0.20) + (100)(100)(0.10) + (100)(200)(0.20) + (250)(0)(0.05)

+ (250)(100)(0.15) + (250)(200)(0.30)

E(XY) = 0 + 1000 + 4000 + 0 + 3750 + 15000 = 23750

So,

Cov(X,Y) = E(XY) - E(X)E(Y) = 23750 - (175)(125) = 23750 - 21875 = 1875

Hence, Cov(x,y) = 1875.